OpenStudy (anonymous):

An armoured vehicle of mass 5000kg is at rest on a frozen lake where friction is practically zero. It fires a shell of mass 10kg with a nuzzle velocity of 800m/s [N]. Let the time take to travel the length of the barrel be "B" seconds. Calculate the acceleration of the shell

OpenStudy (anonymous):

Well, if we assume that the shell accelerates uniformly as it travels through the barrel, we can use \[v_f^2 = v_i^2 + 2a\Delta x\] since v_i = 0, we can solve for a: \[a = \frac{v_f^2}{2\Delta x} \] where delta x is the length of the barrel.

OpenStudy (anonymous):

so my answer could have the variable x in it and it would still be right

OpenStudy (anonymous):

Oh, I'm sorry I misread the question. I should have used the other equation:

OpenStudy (anonymous):

Actually it's not even an equation, it's a definition: \[a = \frac{\Delta v}{t} = \frac{v_f}{B} \]

OpenStudy (anonymous):

i got an answer of \[800m/s/B\]

OpenStudy (anonymous):

does it not give you the time B?

OpenStudy (anonymous):

btw how do you do fractions on this thing and no it doesnt

OpenStudy (anonymous):

Odd. And the syntax is \frac{ numerator }{ denominator }

OpenStudy (anonymous):

inside the \[ braces of course.

OpenStudy (anonymous):

alright thanks. so i guess \[\frac{800m/s}{B}\]

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem